# Hierarchical NeuroSymbolic Approach for Action Quality Assessment

### v1

@@ Abstract
Action quality assessment (AQA) applies computer vision to quantitatively assess the performance or execution of a human action. Current AQA approaches are end-to-end neural models, which lack transparency and tend to be biased because they are trained on subjective human judgements as ground-truth. To address these issues, we introduce a neuro-symbolic paradigm for AQA, which uses neural networks to abstract interpretable symbols from video data and makes quality assessments by applying rules to those symbols. We take diving as the case study. We found that domain experts prefer our system and find it more informative than purely neural approaches to AQA in diving. Our system also achieves state-of-the-art action recognition and temporal segmentation, and automatically generates a detailed report that breaks the dive down into its elements and provides objective scoring with visual evidence. As verified by a group of domain experts, this report may be used to assist judges in scoring, help train judges, and provide feedback to divers. We will open-source all of our annotated training data and code for ease of reproducibility.
@@ 1 Introduction
Analyzing people’s movement and actions and providing feedback has applications ranging from physical rehabilitation [16, 1, 19] to sports coaching and scoring athletic performance [23, 10] to surgery skills assessment [4, 5, 12]. With the 2024 Paris Olympics just around the corner, there is heightened interest in action quality assessment (AQA) in the domain of sports such as diving, gymnastics, and figure skating. These automated systems are generally deep learning-based systems. Such neural approaches analyze videos of sports performance and output a score quantifying how well the athlete performed. These neural models are trained using scores given out by human judges during past competitions.
@@
The issue is that human judging itself is subjective and prone to bias. The bias may take two forms: bias for/against individual athletes; and bias in how different elements of a performative action are weighted in the overall score. An example of the latter in diving is bias towards the last phase of a dive—entry into the water and the amount of splash created. Because of this, the scores given out by the judges may not take into account all the factors/elements of the athletes’ performance. This problem is further aggravated by the fact that performance scores from human judges are never supported by an actual breakdown of the score or an analysis of what athletes did well or not. This ‘black box’ kind of scoring cannot guarantee that: 1) all the factors/elements/efforts from athletes’ performance are taken into account when scoring; & 2) the performance has been objectively scored. This leads to a lack of trust, fairness, and explainability in the scoring. These issues are compounded when current AQA approaches, which are fully neural models [8, 3, 14, 23, 11, 9, 19], are trained using scores given by human judges.
@@ 2 Related Work
Existing approaches treat AQA as a regression problem. Thus far, efforts have focused on predicting action quality scores as closely as possible to those given by human judges. This is achieved through the use of better features [17], multitask learning [18], score distribution learning [21], more data [15], or improved regression techniques [26, 6]. However, the scores given by human judges, which serve as the ’gold standard,’ are inherently biased. For instance, in diving, the last stage of the dive or the splash appears to have the most impact on judges, potentially neglecting other crucial factors of the athlete’s performance. Therefore, at best, these neural approaches may assess in a biased manner.
Moreover, existing approaches being purely neural are ’black box’ in nature—how they arrive at decisions & which factors of the performance are considered, as well as how each factor is scored, remain unknown. This lack of transparency results in a dearth of explainability, fairness, & trustworthiness, making it uncertain whether all the efforts from performers are genuinely rewarded. Ideally, athletes who have sacrificed a significant part of their lives should receive recognition for all their efforts. In contrast, we propose a new paradigm to AQA—a Neuro-Symbolic approach. Our approach leverages modern deep neural networks to extract useful information from performance videos into symbols. It then applies logic to these symbols to make decisions and evaluate performance without relying on biased ground-truths, but rather utilizes domain knowledge for objective scoring. Our system conducts a much more comprehensive and in-depth evaluation, programmatically generating an extremely detailed performance evaluation report. This makes the evaluation process transparent, fair, explainable, and trustworthy.
Neurosymbolic AI which combines the power of neural models and symbolic systems has been gaining increasing interest [2, 25, 20], but has not been explored in AQA setting. In this work, we develop a hierarchical neurosymbolic approach to AQA.
@@ 3 Hierarchical Neuro-Symbolic Approach
Platform. The location of the platform, especially the position of its edge facing the pool, is crucial to determine when the diver leaves the platform, thus starting their dive. The platform location is also important to assess how close the diver comes to its edge, which is relevant to scoring.
Splash at entry into the pool is a conspicuous visual feature of a dive. The size of the splash is an important element in traditional scoring of dives. A large splash mars the end of a dive and also likely indicates a flaw in form at water entry.
Pose of the diver in video frames is crucial for dive assessment. We gather 2D pose data, detailing body part locations like the head, thorax, pelvis, etc. Analyzing this data frame by frame allows identification of sub-actions (e.g., somersaults, twists) and quality assessment. Detailed pose data enables objective metrics like water entry angle.
@@ 3.1 Neural Action-Context Parser [Uncaptioned image]
We implement the Neural Action-Context Parser as a collection of a dedicated: 1) splash detector; 2) platform detector, and 3) human pose estimator. These object detectors and pose estimators first extract relevant information from each frame individually. For instance, we run a splash detector on all frames; whenever a splash is detected, it is segmented out. Object segmentation provides access to its properties, such as 1) size, 2) shape, and 3) position within the scene. In cases where no splash is detected, all properties are filled with null values. Similarly, pose estimators record 2D positions of all major joints of the diver. Direct information from these detectors forms the primary set of symbols. We further derive relations between the primary symbols, such as: 1) the angles made by the bones at joints, and 2) the distance between the diver and the platform. These form a secondary set of symbols. Additional information is provided in Table 1.
@@ 3.2.3 Fine-grained Dive (Action) Quality Assessment
Distance-from-platform of a diver is crucial for safety and aesthetic reasons. Divers risk injury or even death by hitting the concrete platform by being too close to it, leading to penalties. Conversely, being too far from the platform results in an unaesthetic trajectory and is also penalized.
Feet-apart. A streamlined and graceful form throughout a dive is considered ideal. Taking a cue from gymnastics, one aspect of an ideal diving form is maintaining one’s feet together. A diver is penalized for having their feet apart.
Somersault tightness. During a somersault, it is desirable for a diver to have as tight a tuck or pike position as possible. A very tight somersault position allows a diver to spin faster and is considered ideal.
Over/under-rotation. Each particular dive has a specified number of rotations of somersaults to be performed. The number of rotations is in half-rotation increments. A diver preferably performs the exact number of rotations (to the degree) that are specified for the dive–no more and no less. Over rotation and under rotation are penalized.
Splash size. Ideally, a diver has little or no splash upon entry into the water. Little or no splash after a dive from 10 meters high is difficult to achieve and is indicative of very good form upon entry.
@@ 4.3.2 Expert Feedback
To validate the practical usefulness of our system in relation to diving competitions & training, we presented our system & its automatically generated reports (explained in Approach) to the Team USA 2016 Olympic diving coach, Oleg Andriyuk, & Chair of the USA Diving Judges Commission, Amy Kwan. Both of them have generously offered valuable feedback for our system.
Expert Opinion 1: Coach Andriyuk.
According to Andriyuk, human judges emphasize overall impression when scoring a dive. This is because they cannot always see all of the diver’s mistakes in the short amount of time watching the dive so as to come up with a score based on specific errors. Even if judges try to be fair to all aspects and phases of a dive, as humans, it is almost impossible to see the dive perfectly at all times. Moreover, judging based on overall impression is inherently subjective and vulnerable to bias. Andriyuk believes that with the support of our NS-AQA system, scoring in diving can be “more precise and have less errors.” Andriyuk believes that this system, with its generated detailed report, would be helpful in a number of ways. In particular, it would be useful as a tool to (1) teach judges how to score, (2) catch any errors that a human judge may miss, (3) settle disagreements between judges, and (4) encourage safety by penalizing dangerous actions (e.g. getting too close to hitting the board) that lots of human judges overlook. See full opinion in supplementary.
Expert Opinion 2: Judge Kwan.
Kwan envisioned our system as an educational tool for teaching judges, coaches, and divers how to break down the dive into all of its components. “We’re humans, we can’t get it all right,” Kwan says, especially since judges only get 3 seconds to score a dive. Kwan particularly emphasized our system’s potential to improve diving safety due to its automated “distance from board” measurement. In diving, hitting the platform can be the difference between life and death, so it is very important to prioritize the diver’s safety. She is “very excited for this system to be able to make that call for us,” as it moves the villain-role from the human judges to the AI to make those difficult decisions. See full opinion in supplementary.
@@ 4.3.3 Comparison to Neural Approach C3D-MTL
Key Frames of Dive
Our Output (Summarized, not full)
Neural Model Score Comparison
@@
Overall Score: 5.7
Feet Apart: 2.2
Height off Platform: 2.7
Distance from Platform: Too Far
Somersault Tightness: 4.1
Knee Straightness: 7.8
Twist Straightness: 5.5
Verticalness: 6.3
Body Straightness during Entry: 8.2
Splash: 9.5
Overall Score: 9.7 Note that, existing neural models only output a final score, unlike our model, which outputs a detailed report along with the final score.
The score outputted by C3D-MTL was significantly higher than our system. We see by looking at the dive’s key frames that the start/takeoff and flight was loose and sloppy, but the entry was extremely good. This suggests that the neural model was only factoring the entry into the overall score instead of also accounting for the beginning parts of the dive.
@@
Overall Score: 5.4
Feet Apart: 3.9
Height off Platform: 0.1
Distance from Platform: Good
Somersault Tightness: 0.9
Knee Straightness: 7.8
Twist Straightness: 6.0
Verticalness: 4.8
Body Straightness during Entry: 9.5
Splash: 5.9
Overall Score: 7.8
The score outputted by C3D-MTL was much higher than our system. We see by looking at the dive’s key frames that the height off the platform was very low and the pike in the somersault was not tight. However, the entry was with little splash. This suggests that the neural model was mostly factoring the splash into the overall score instead of also accounting for factors like height and position tightness.
@@
Overall Score: 6.4
Feet Apart: 8.8
Height off Platform: 8.5
Distance from Platform: Good
Somersault Tightness: 6.0
Knee Straightness: N/A
Twist Straightness: N/A
Verticalness: 1.3
Body Straightness during Entry: 9.8
Splash: 0.0
Overall Score: 3.4
The score outputted by C3D-MTL was much lower than our system. Every aspect of the dive including the height and position tightness was very good except for a slight mistiming when the diver came out of the dive, resulting in over-rotation. The diver should be rewarded for the other components of the dive they did well, instead of only penalized for their poor entry.
Table 9: Qualitative Comparison of Our Neural-Symbolic System to the Neural Model C3D-MTL. As taught in the official USA Diving Judge’s Course, “Dives should be judged as a whole, not emphasizing any single element. This approach in diving is especially true when it comes to the entry. It is easy to forgive earlier flaws if the dive has a good entry.” [7] In all three examples, we see that the neural model emphasizes the entry of the dive far more than any other aspect, while our system more fairly accounts for each aspect.
@@ 5 Discussion
We propose a neuro-symbolic (NS) paradigm for AQA, & implement an NS-AQA system specifically for scoring competitive diving. Our NS system achieves SOTA results for fine-grained action recognition & temporal segmentation, & automatically generates a detailed & comprehensive report scoring the performance of the dive. Our system was validated by a group of domain experts including Olympic diving coach & the USA Diving Judges Commission Chair. Our approach ensures fairness and transparency, minimizing biases that often plague traditional assessment methods. Unlike traditional computer vision-based systems & human judges that offer single scores without explanation, our NS approach provides a breakdown of assessments, empowering athletes & coaches with actionable insights.
Extension to other actions and skills assessment. Although this paper is geared towards diving, our NS approach is general in nature, readily extensible to other sports like the vault in gymnastics & figure skating. Domain expertise would be required to formulate rules. Our approach can be extended to other critical domains like surgery skills assessment as well. In surgery skills assessment, for example, microprograms to detect bleeding (analogous to splash in our case), respect for tissue, etc. could be developed. Final rating would be an aggregation of all such factors. We will release our codebase and data, hoping to inspire its application across diverse domains.

### v2

# Hierarchical NeuroSymbolic Approach for Comprehensive and Explainable Action Quality Assessment

@@ Abstract
Action quality assessment (AQA) applies computer vision to quantitatively assess the performance or execution of a human action. Current AQA approaches are end-to-end neural models, which lack transparency and tend to be biased because they are trained on subjective human judgements as ground-truth. To address these issues, we introduce a neuro-symbolic paradigm for AQA, which uses neural networks to abstract interpretable symbols from video data and makes quality assessments by applying rules to those symbols. We take diving as the case study. We found that domain experts prefer our system and find it more informative than purely neural approaches to AQA in diving. Our system also achieves state-of-the-art action recognition and temporal segmentation, and automatically generates a detailed report that breaks the dive down into its elements and provides objective scoring with visual evidence. As verified by a group of domain experts, this report may be used to assist judges in scoring, help train judges, and provide feedback to divers. Annotated training data and code: https://github.com/laurenok24/NSAQA.
@@ 1 Introduction
Analyzing people’s movement and actions and providing feedback has applications ranging from physical rehabilitation [19, 1, 22] to sports coaching and scoring athletic performance [28, 11, 15] to surgery skills assessment [5, 6, 14]. With the 2024 Paris Olympics just around the corner, there is heightened interest in action quality assessment (AQA) in the domain of sports such as diving, gymnastics, and figure skating. These automated systems are generally deep learning-based systems. Such neural approaches analyze videos of sports performance and output a score quantifying how well the athlete performed. These neural models are trained using scores given out by human judges during past competitions.
@@
The issue is that human judging itself is subjective and prone to bias. The bias may take two forms: bias for/against individual athletes; and bias in how different elements of a performative action are weighted in the overall score. An example of the latter in diving is bias towards the last phase of a dive—entry into the water and the amount of splash created. Because of this, the scores given out by the judges may not take into account all the factors/elements of the athletes’ performance. This problem is further aggravated by the fact that performance scores from human judges are never supported by an actual breakdown of the score or an analysis of what athletes did well or not. This ‘black box’ kind of scoring cannot guarantee that: 1) all the factors/elements/efforts from athletes’ performance are taken into account when scoring; & 2) the performance has been objectively scored. This leads to a lack of trust, fairness, and explainability in the scoring. These issues are compounded when current AQA approaches, which are fully neural models [9, 4, 17, 28, 12, 10, 22, 2, 33], are trained using scores given by human judges.
@@ 2 Related Work
Existing approaches treat AQA as a regression problem. Thus far, efforts have focused on predicting action quality scores as closely as possible to those given by human judges. This is achieved through the use of better features [20], multitask learning [21], score distribution learning [25], more data [18], or improved regression techniques [31, 7]. However, the scores given by human judges, which serve as the ‘gold standard,’ are inherently biased. For instance, in diving, the last stage of the dive or the splash appears to have the most impact on judges, potentially neglecting other crucial factors of the athlete’s performance. Therefore, at best, these neural approaches may assess in a biased manner.
Moreover, existing approaches being purely neural are ’black box’ in nature—how they arrive at decisions & which factors of the performance are considered, as well as how each factor is scored, remain unknown. This lack of transparency results in a dearth of explainability, fairness, & trustworthiness, making it uncertain whether all the efforts from performers are genuinely rewarded. Ideally, athletes who have sacrificed a significant part of their lives should receive recognition for all their efforts. In contrast, we propose a new paradigm to AQA—a Neuro-Symbolic approach. Our approach leverages modern deep neural networks to extract useful information from performance videos into symbols. It then applies logic to these symbols to make decisions and evaluate performance without relying on biased ground-truths, but rather utilizes domain knowledge for objective scoring. Our system conducts a much more comprehensive and in-depth evaluation, programmatically generating an extremely detailed performance evaluation report. This makes the evaluation process transparent, fair, explainable, and trustworthy. A concurrent work [32] also recognizes the value of detailed action evaluation; however, it relies on large specialized datasets and, being a purely neural approach, still suffers from previously mentioned issues. Additionally, our reports are more comprehensive and detailed, complete with visual evidence.
Neurosymbolic AI which combines the power of neural models and symbolic systems has been gaining increasing interest [3, 30, 23, 13], but has not been explored in AQA setting. In this work, we develop a hierarchical neurosymbolic approach to AQA.
@@ 3 Hierarchical Neuro-Symbolic Approach
Platform. The location of the platform, especially the position of its edge facing the pool, is crucial to determine when the diver leaves the platform, thus starting their dive. The platform location is also important to assess how close the diver comes to its edge, which is relevant to scoring. [Uncaptioned image]
Splash at entry into the pool is a conspicuous visual feature of a dive. The size of the splash is an important element in traditional scoring of dives. A large splash mars the end of a dive and also likely indicates a flaw in form at water entry. [Uncaptioned image]
Pose of the diver in video frames is crucial for dive assessment. We gather 2D pose data, detailing body part locations like the head, thorax, pelvis, etc. Analyzing this data frame by frame allows identification of sub-actions (e.g., somersaults, twists) and quality assessment. Detailed pose data enables objective metrics like water entry angle. [Uncaptioned image]
@@ 3.1 Neural Action-Context Parser [Uncaptioned image]
We implement the Neural Action-Context Parser as a collection of a dedicated: 1) splash detector; 2) platform detector, and 3) human pose estimator. These object detectors [27] and pose estimators [24] first extract relevant information from each frame individually. For instance, we run a splash detector on all frames; whenever a splash is detected, it is segmented out. Object segmentation provides access to its properties, such as 1) size, 2) shape, and 3) position within the scene. In cases where no splash is detected, all properties are filled with null values. Similarly, pose estimators record 2D positions of all major joints of the diver. Direct information from these detectors forms the primary set of symbols. We further derive relations between the primary symbols, such as: 1) the angles made by the bones at joints, and 2) the distance between the diver and the platform. These form a secondary set of symbols. Additional information is provided in Table 1.
@@ 3.2.3 Fine-grained Dive (Action) Quality Assessment
Distance-from-platform of a diver is crucial for safety and aesthetic reasons. Divers risk injury or even death by hitting the concrete platform by being too close to it, leading to penalties. Conversely, being too far from the platform results in an unaesthetic trajectory and is also penalized. [Uncaptioned image]
Feet-apart. A streamlined and graceful form throughout a dive is considered ideal. Taking a cue from gymnastics, one aspect of an ideal diving form is maintaining one’s feet together. A diver is penalized for having their feet apart. [Uncaptioned image]
Somersault tightness. During a somersault, it is desirable for a diver to have as tight a tuck or pike position as possible. A very tight somersault position allows a diver to spin faster and is considered ideal. [Uncaptioned image]
Over/under-rotation. Each particular dive has a specified number of rotations of somersaults to be performed. The number of rotations is in half-rotation increments. A diver preferably performs the exact number of rotations (to the degree) that are specified for the dive–no more and no less. Over rotation and under rotation are penalized. [Uncaptioned image]
Splash size. Ideally, a diver has little or no splash upon entry into the water. Little or no splash after a dive from 10 meters high is difficult to achieve and is indicative of very good form upon entry. [Uncaptioned image]
@@ 4.3.2 Expert Feedback
To validate the practical usefulness of our system in relation to diving competitions & training, we presented our system & its automatically generated reports (explained in Approach) to an Olympic diving coach (referred to as Coach A for anonymity), & a certified judge on the USA Diving Judges Commission (referred to as Judge B). Both of them have generously offered valuable feedback for our system.
Expert Opinion 1: Coach A.
According to Coach A, human judges emphasize overall impression when scoring a dive. This is because they cannot always see all of the diver’s mistakes in the short amount of time watching the dive so as to come up with a score based on specific errors. Even if judges try to be fair to all aspects and phases of a dive, as humans, it is almost impossible to see the dive perfectly at all times. Moreover, judging based on overall impression is inherently subjective and vulnerable to bias. Coach A believes that with the support of our NS-AQA system, scoring in diving can be “more precise and have less errors.” Coach A believes that this system, with its generated detailed report, would be helpful in a number of ways. In particular, it would be useful as a tool to (1) teach judges how to score, (2) catch any errors that a human judge may miss, (3) settle disagreements between judges, and (4) encourage safety by penalizing dangerous actions (e.g. getting too close to hitting the board) that lots of human judges overlook. See full opinion in supplementary.
Expert Opinion 2: Judge B.
Judge B envisioned our system as an educational tool for teaching judges, coaches, and divers how to break down the dive into all of its components. “We’re humans, we can’t get it all right,” Judge B says, especially since judges only get 3 seconds to score a dive. Judge B particularly emphasized our system’s potential to improve diving safety due to its automated “distance from board” measurement. In diving, hitting the platform can be the difference between life and death, so it is very important to prioritize the diver’s safety. Judge B is “very excited for this system to be able to make that call for us,” as it moves the villain-role from the human judges to the AI to make those difficult decisions. See full opinion in supplementary.
@@
Key Frames of Dive Our Output (Summarized, not full) Neural Model Score Comparison
@@
Overall Score: 5.7 Feet Apart: 2.2 Height off Platform: 2.7 Distance from Platform: Too Far Somersault Tightness: 4.1 Knee Straightness: 7.8 Twist Straightness: 5.5 Verticalness: 6.3 Body Straightness during Entry: 8.2 Splash: 9.5 Overall Score: 9.7 Note that, existing neural models only output a final score, unlike our model, which outputs a detailed report along with the final score. The score outputted by C3D-MTL was significantly higher than our system. We see by looking at the dive’s key frames that the start/takeoff and flight was loose and sloppy, but the entry was extremely good. This suggests that the neural model was only factoring the entry into the overall score instead of also accounting for the beginning parts of the dive.
@@
Overall Score: 5.4 Feet Apart: 3.9 Height off Platform: 0.1 Distance from Platform: Good Somersault Tightness: 0.9 Knee Straightness: 7.8 Twist Straightness: 6.0 Verticalness: 4.8 Body Straightness during Entry: 9.5 Splash: 5.9 Overall Score: 7.8 The score outputted by C3D-MTL was much higher than our system. We see by looking at the dive’s key frames that the height off the platform was very low and the pike in the somersault was not tight. However, the entry was with little splash. This suggests that the neural model was mostly factoring the splash into the overall score instead of also accounting for factors like height and position tightness.
@@
Overall Score: 6.4 Feet Apart: 8.8 Height off Platform: 8.5 Distance from Platform: Good Somersault Tightness: 6.0 Knee Straightness: N/A Twist Straightness: N/A Verticalness: 1.3 Body Straightness during Entry: 9.8 Splash: 0.0 Overall Score: 3.4 The score outputted by C3D-MTL was much lower than our system. Every aspect of the dive including the height and position tightness was very good except for a slight mistiming when the diver came out of the dive, resulting in over-rotation. The diver should be rewarded for the other components of the dive they did well, instead of only penalized for their poor entry.
Table 9: Qualitative comparison of our Neuro-Symbolic system to the purely Neural model C3D-MTL. As taught in the official USA Diving Judge’s Course, “Dives should be judged as a whole, not emphasizing any single element. This approach in diving is especially true when it comes to the entry. It is easy to forgive earlier flaws if the dive has a good entry.” [8] In all three examples, we see that the neural model emphasizes the entry of the dive far more than any other aspect, while our system more fairly accounts for each aspect.
4.3.3 Comparison to Neural Approach C3D-MTL
@@ 5 Discussion
We propose a neuro-symbolic (NS) paradigm for AQA, & implement an NS-AQA system specifically for scoring competitive diving. Our NS system achieves SOTA results for fine-grained action recognition & temporal segmentation, & automatically generates a detailed & comprehensive report scoring the performance of the dive. Our system was validated by domain experts including an Olympic diving coach & a certified judge on the Judges Commission. Our approach ensures fairness and transparency, minimizing biases that often plague traditional assessment methods. Unlike traditional computer vision-based systems & human judges that offer single scores without explanation, our NS approach provides a breakdown of assessments, empowering athletes & coaches with actionable insights.
Extension to other actions and skills assessment. Although this paper is geared towards diving, our NS approach is general in nature, readily extensible to other sports like the vault in gymnastics & figure skating. Domain expertise would be required to formulate rules. Our approach can be extended to other critical domains like surgery skills assessment as well. In surgery skills assessment, for example, microprograms to detect bleeding (analogous to splash in our case), respect for tissue, etc. could be developed. Final rating would be an aggregation of all such factors. We hope to inspire this work’s application across diverse domains.
Acknowledgement. LO was supported by Princeton University.